OpenStudy (anonymous):

You spin a spinner that has eight equal-sized sections numbered one to eight. Find P(perfect square or odd). 1/8 3/8 5/8 6/8

jimthompson5910 (jim_thompson5910):

In the numbers from 1 to 8: 1,2,3,4,5,6,7,8 How many perfect squares are there?

OpenStudy (kropot72):

Is 7 a perfect square?

OpenStudy (anonymous):

i think so

OpenStudy (kropot72):

\[\sqrt{7}=?\]

OpenStudy (anonymous):

1/8?

OpenStudy (kropot72):

\[\frac{1}{8}\times \frac{1}{8}=\frac{1}{64}\]

jimthompson5910 (jim_thompson5910):

There are 2 perfect squares: 1, 4 and 4 odd numbers: 1,3,5,7 So there are 6 numbers total out of 8 giving us the probability 6/8